How many pounds of cashews selling for $7.40 per pound must be mixed with 6 pounds of peanuts selling for $2.60 per pound to create a mixture that sells for $5.80 per pound?
Answer by jorel1380(3719)

7.40x+6(2.60)=5.80(x+6)
7.40x+15.6=5.80x+34.8
1.6x=19.2
x=12
You must add 12 pounds of cashews to the peanuts to get a mixture that sells for $5.80 a pound.